# Show and Tell: Visually Explainable Deep Neural Nets via Spatially-Aware Concept Bottleneck Models

### v3

@@ 3.2 Local image-concept similarities
Vision-language models such as CLIP have been widely used for obtaining global image descriptions, as in non-spatial CBMs [36, 22]. Here, we aim to extend this approach to locally describe different image regions using visual concepts. Inspired by [27], we leverage CLIP’s visual prompting property by drawing a red circle around specific image regions, which enables CLIP to focus on these areas while preserving global context. We apply this property to our training set as illustrated in Figure 2(b). Formally, let $x_{n}\in\mathcal{X}$ denote a training image with spatial dimensions ${H\times W}$. We create a uniform grid of $\tilde{H}\times\tilde{W}$ locations in the image with integer strides ${d_{H}}$ and ${d_{W}}$, i.e., $d_{H}=\lfloor\frac{H}{\tilde{H}-1}\rfloor$ and $d_{W}=\lfloor\frac{W}{\tilde{W}-1}\rfloor$. For each $x_{n}$, a set of $\tilde{H}\cdot\tilde{W}$ augmented images is then obtained by drawing a red circle with radius $r$ around each location in the grid. We denote by $x_{n}^{(h,w)}$ the image $x_{n}$ with a red circle located at ${(h,w)}$. Next, we compute a local similarity score between a visual concept $t_{m}\in\mathcal{T}$ and the image at location ${(h,w)}$ as follows: $P[n,m,h,w]=\frac{I_{n}^{(h,w)}\cdot T_{m}}{\lVert I_{n}^{(h,w)}\rVert\lVert T_{m}\rVert}$, where $I_{n}^{(h,w)}=E_{I}(x_{n}^{(h,w)})$ and $T_{m}=E_{T}(t_{m})$ denote the CLIP embeddings of the augmented image and the concept, respectively. As demonstrated in [27], this score measures how well the concept $t_{m}$ describes the image region highlighted by the red circle. By computing this score for every spatial location $(h,w)$ in the grid, we obtain a concept similarity map for the entire image $x_{n}$, as shown in Figure 2(b). A complete spatial concept similarity matrix $P$ is then constructed by calculating these local similarity scores for all concepts $t_{m}\in\mathcal{T}$ and all images $x_{n}\in\mathcal{X}$ in the training set. This matrix is computed once, prior to training, and is later used for learning the spatially-aware concept bottleneck layer. We note that the grid resolution $\tilde{H}\times\tilde{W}$ and the circle radius $r$ are hyper-parameters, where $\tilde{H}$ and $\tilde{W}$ control the coarseness of the concept similarity map, and $r$ defines the receptive field around each location. We optimize these hyper-parameters per-dataset (see supplementary).
@@ 3.3 Training the concept-bottleneck layer
[…]
@@ 3.4 Training the final classification layer
[…]
@@ 5 Conclusions
Additionally, we introduced interactive capabilities for model exploration and debugging, demonstrating their effectiveness in diagnosing and correcting model errors. We believe that such features are particularly valuable for high-stakes applications like medical imaging and autonomous driving. By providing expert practitioners with intuitive tools to understand and adjust model decisions, our approach can boost confidence and support safer deployment in these critical fields. Looking ahead, as new VLMs are developed across various domains, our findings could help inform the design of more powerful interpretability tools for a broad spectrum of AI applications. We plan to explore these directions in future work.
@@ A.1 Local image-concept similarities
Input: (i) training images $\{x_{n}\}^{N}_{n=1}\in\mathbb{R}^{3\times H\times W}$; (ii) concept list $\{t_{m}\}^{M}_{m=1}$; (iii) CLIP’s image encoder $E_{I}$ and text encoder $E_{T}$; (iv) circle radius $r$, grid dimensions $(\tilde{H},\tilde{W})$.
@@
$T\leftarrow E_{T}(\{t_{m}\}_{m=1}^{M})$triangleright encode all concepts once
for $n\leftarrow 0$ to $N-1$ do $\triangleright$ iterate over training images
@@
     $I_{n}\leftarrow E_{I}(X_{n}^{aug})$triangleright encode entire augmented batch

### v4

@@ 3.2 Local image-concept similarities
Vision-language models such as CLIP have been widely used for obtaining global image descriptions, as in non-spatial CBMs [36, 22]. Here, we aim to extend this approach to locally describe different image regions using visual concepts. Inspired by [27], we leverage CLIP’s visual prompting property by drawing a red circle around specific image regions, which enables CLIP to focus on these areas while preserving global context. We apply this property to our training set as illustrated in Figure 2(b). Formally, let $x_{n}\in\mathcal{X}$ denote a training image with spatial dimensions $H\times W$}. We create a uniform grid of $\tilde{H}\times\tilde{W}$ locations in the image with integer strides $d_{H}$} and $d_{W}$}, i.e., $d_{H}=\lfloor\frac{H}{\tilde{H}-1}\rfloor$ and $d_{W}=\lfloor\frac{W}{\tilde{W}-1}\rfloor$. For each $x_{n}$, a set of $\tilde{H}\cdot\tilde{W}$ augmented images is then obtained by drawing a red circle with radius $r$ around each location in the grid. We denote by $x_{n}^{(h,w)}$ the image $x_{n}$ with a red circle located at (h,w){(h,w)}. Next, we compute a local similarity score between a visual concept $t_{m}\in\mathcal{T}$ and the image at location (h,w){(h,w)} as follows: $P[n,m,h,w]=\frac{I_{n}^{(h,w)}\cdot T_{m}}{\lVert I_{n}^{(h,w)}\rVert\lVert T_{m}\rVert}$, where $I_{n}^{(h,w)}=E_{I}(x_{n}^{(h,w)})$ and $T_{m}=E_{T}(t_{m})$ denote the CLIP embeddings of the augmented image and the concept, respectively. As demonstrated in [27], this score measures how well the concept $t_{m}$ describes the image region highlighted by the red circle. By computing this score for every spatial location $(h,w)$ in the grid, we obtain a concept similarity map for the entire image $x_{n}$, as shown in Figure 2(b). A complete spatial concept similarity matrix $P$ is then constructed by calculating these local similarity scores for all concepts $t_{m}\in\mathcal{T}$ and all images $x_{n}\in\mathcal{X}$ in the training set. This matrix is computed once, prior to training, and is later used for learning the spatially-aware concept bottleneck layer. We note that the grid resolution $\tilde{H}\times\tilde{W}$ and the circle radius $r$ are hyper-parameters, where $\tilde{H}$ and $\tilde{W}$ control the coarseness of the concept similarity map, and $r$ defines the receptive field around each location. We optimize these hyper-parameters per-dataset (see supplementary).
@@ 3.3 Training the concept-bottleneck layer
$\mathcal{L}_{CBL}=-\underset{m=1}{\overset{M}{\sum}}\underset{h,w}{\sum}sim\left(q[m,h,w],p[m,h,w]\right)$ (1)
@@ 3.4 Training the final classification layer
$\underset{n=1}{\overset{N}{\sum}}\mathcal{L}_{ce}\left(Wc^{\ast}(x_{n})+b,y_{n}\right)+\lambda\mathcal{R}(W)$ (2)
@@ 5 Conclusions
Additionally, we introduced interactive capabilities for model exploration and debugging, demonstrating their effectiveness in diagnosing and correcting model errors. We believe that such features are particularly valuable for high-stakes applications like medical imaging and autonomous driving. By providing expert practitioners with intuitive tools to understand and adjust model decisions, our approach can boost confidence and support safer deployment in these critical fields. As new VLMs are developed across various domains, our findings could help the design of more powerful interpretability tools for a broad spectrum of applications. We plan to explore these directions in future work.
Acknowledgments
This work was partially supported by the Israel Ministry of Science and Technology (T.R.R MOST 0007311/1001817977).
@@ A.1 Local image-concept similarities
[…]
@@
$T\leftarrow E_{T}(\{t_{m}\}_{m=1}^{M})$ $\triangleright$ encode all concepts once
for $n\leftarrow 0$ to N1N-1 do $\triangleright$ iterate over training images
@@
  $I_{n}\leftarrow E_{I}(X_{n}^{aug})$ $\triangleright$ encode entire augmented batch